how many grams of NaOH are produced when 3.20 grams of O2 is formed?
1
Expert's answer
2012-05-17T09:58:02-0400
NaOH and O2 are formed when Na2O2 reacted with water by reaction:
2Na2O2 +
2H2O = 4NaOH + O2
3.2 g of O2 corresponds to 0.1 mol of O2
Thereby, the
quantity of NaOH is 4*0.1=0.4 mol
The mass of NaOH is m(NaOH) =
M(NaOH)*n(NaOH) = 40g/mol * 0.4 mol = 16 g
Answer: 16 grams of NaOH are
produced when 3.20 grams of O2 is formed.